OpenStudy (anonymous):

150=50e^(7x), solve for x.

OpenStudy (anonymous):

divide both sides by 50, and then ln both sides.

OpenStudy (anonymous):

divide by 50 get \[3=e^{7x}\] then rewrite in equivalent logarithmic form as \[\ln(3)=7x\] then solve for \(x\)

OpenStudy (anonymous):

you'd get e^(7x)=30

OpenStudy (anonymous):

which really just means divide by \(7\) if you want a decimal answer use a calculator to compute \[\frac{\ln(3)}{7}\]

OpenStudy (anonymous):

\[150\div 50=3\]

OpenStudy (anonymous):

thank you so much @satellite73

OpenStudy (anonymous):

yw